OpenStudy (anonymous):

Anyone in Connections taking algebra 2 B?

OpenStudy (anonymous):

Need help with the get the picture portfolio

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have the get the picture portfolio done?

OpenStudy (anonymous):

dno

OpenStudy (anonymous):

did you do this?

OpenStudy (anonymous):

i can't open i

OpenStudy (solomonzelman):

I am in Connections, but not in so called course. I would suggest asking a question directly, unless this is a classmate search - which is more appropriate for a Feedback section.

OpenStudy (solomonzelman):

srry nvm

OpenStudy (anonymous):

@doc.brown pls tell me your in connections?

OpenStudy (doc.brown):

Just write the question down, I will help.

OpenStudy (anonymous):

• Measure the lengths of the arrows shown at the right. What is the relationship between these lengths? How does this relate to your research on perspective? • Trace the four arrows at the right, moving the paper to the left after tracing the longest arrow so that it is further away from the others than it is now. What do you notice? • Make a simple drawing of three or more similar objects whose lengths can be represented by an arithmetic sequence. Write the corresponding arithmetic sequence, and a recursive or explicit formula for that sequence.

OpenStudy (anonymous):

the firs arrow is 2in, the 2nd one is 1 1/2, the 3rd is 1 1/4, and the forth is 1 in

OpenStudy (doc.brown):

\[a_1=2\]\[a_2=1\frac{1}{2}\]\[a_3=1\frac{1}{4}\]\[a_4=1\]

OpenStudy (anonymous):

if it says that im not here, it's because of the lag lol sorry

OpenStudy (doc.brown):

\[a_1=2\]\[a_2=1\frac{1}{2}=\frac{3}{2}\]\[a_3=1\frac{1}{4}=\frac{5}{4}\]\[a_4=1\]

OpenStudy (doc.brown):

\[a_1=2=\frac{8}{4}\]\[a_2=1\frac{1}{2}=\frac{3}{2}=\frac{6}{4}\]\[a_3=1\frac{1}{4}=\frac{5}{4}\]\[a_4=1=\frac{4}{4}\]

OpenStudy (doc.brown):

See the relationship now? 4:5:6:8

OpenStudy (anonymous):

Yes!

OpenStudy (doc.brown):

I imagine the arrows are in a 3D landscape, you teacher is likely looking for a small rant about how things further away appear smaller, depending on your point of view? I don't know your research.

OpenStudy (anonymous):

okay i was going to say that for the second part

OpenStudy (anonymous):

now the third part has me at a lost

OpenStudy (anonymous):

this is the picture i want to use

OpenStudy (doc.brown):

Just make up some number, like 2 or 3/7 or something. Draw one length 1 and another some multiple of your made up number. 2x or 3x/7

OpenStudy (doc.brown):

Ok, you add three to each pile. Recursive. n+3

OpenStudy (doc.brown):

\[a_1=2\]\[a_n=a_{n-1}+3\]

OpenStudy (anonymous):

okay so my arithmetic sequence would be 2:5:8:11:14 etc. and the recursive is what you just wrote?

OpenStudy (doc.brown):

Yup, double check to make sure what I wrote makes sense. start with \(a_1\), which is 2 then make \(n=2\) \(a_2=a_1+3=5\)\[a_3=a_2+3=8\]

OpenStudy (doc.brown):

keep going for n=3, n=4, n=5, etc Look good?

OpenStudy (anonymous):

yup thanks! can you help with another one?

OpenStudy (anonymous):

file:///C:/Users/jennaci4/Pictures/621504-1132012-110216-AM-1919289153.pdf

OpenStudy (doc.brown):

Sure, if I get a medal here, best response button. Then close the question and post a new one. I'll come find you.

OpenStudy (anonymous):

okay thanks ill tag you!